CHEM 1

Hydrazine (N2H4) and dinitrogen tetroxide ignite on contact to form nitrogen gas and water vapor. How many grams of nitrogen gas form when 1.00 X 102 grams of N2H4 and2.00 X 102 grams of N2O4 are mixed? Which reactant is the limiting reactant? Please explain the steps.
Answer #1

2N2H4 + N2O4 ===> 4H2O + 3N2

100/32/2 = 1.5625 moles of N2O4 can be reacted by N2H4

200/92 = 2.174 moles of N2O4 are present

Thus N2H4 is limiting

Grams of N2 = 100/32 moles of N2H4 * (3molesN2/2 moles N2H4) * 28grams N2/mole = 131 grams of N2 can be produced